OpenStudy (anonymous):

Graph the following function by starting with the graph of y = x^2 and using transformations (shifting, compressing, stretching, and/or reflection). f(x) = -2x^2 + 6x + 2

OpenStudy (anonymous):

I already got the first step down of putting it like: -2x^2 + 6x + 2 = -2x^2 + 6x + 9 - 9 + 2

jigglypuff314 (jigglypuff314):

you should first factor out the greatest common factor, or the way you're doing it won't work...

OpenStudy (anonymous):

greatest common factor is x?

jigglypuff314 (jigglypuff314):

f(x) = -2x^2 + 6x + 2 = -2(x^2 + 3x) + 2 = -2(x^2 + 3x + 2.25 - 2.25) + 2 = -2(x^2 + 3x + 2.25) - 4.5 + 2 = -2(x + 1.5)^2 - 2.5

OpenStudy (anonymous):

how do i graph that?

jigglypuff314 (jigglypuff314):

-2(x + 1.5)^2 - 2.5 ^ is how much it is stretched ^ how much it is shifted horizontally ^ how much it is shifted vertically so flip parent graph upside down because of the negative in the -2 then stretch vertically by 2 then move it 1.5 to the left and then move it 2.5 down.... do you see where I'm getting my numbers?

OpenStudy (anonymous):

yeah i get it! thanks for the help-